OpenStudy (anonymous):

A piece of art is in the shape of an equilateral triangle with sides of 6m. Find the area of the piece of art.

OpenStudy (johnweldon1993):

you know the area of a triangle is \[\frac{ 1 }{ 2 } b h\]  as you can see...you have 2 right triangles the hypotenuse is 6...and the bottom side is 6/2 = 3 so now you have 1 side...the hypotenuse...find the other side *which will be your height* using the Pythagorean theorem 3² + b² = 6²

OpenStudy (anonymous):

thats the answer?

OpenStudy (johnweldon1993):

no ..what you get from b = √(6² - 3²) will be your height Then you plug THAT into that other equation along with the base length (6) \[\frac{ 1 }{ 2 } (6) (height)\]

OpenStudy (anonymous):

I lost my calc...:(

OpenStudy (johnweldon1993):

well lets see 36 - 9 = 27 √ 27 = 5.19615242270 that is your height \[\frac{ 1 }{ 2 } (6) (5.19615242270)\] and that comes out to 15.59 would be the area

OpenStudy (anonymous):

thanks :)

OpenStudy (johnweldon1993):

anytime :)

OpenStudy (anonymous):

Do go to CA?

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

No I do not sorry, I'm actually in University :)